Can you please help me solve this question? One computer printer can print a company's mailing labels in 40 minutes. A second printer would take 60 minutes to print the labels. How long would it take the two printers, operating together, to print the labels? Thank you.
Answer by ptaylor(2198)

Let x= time it takes both computers working together
Then both computers print at the rate of 1/x mailing labels per min
First computer prints at the rate of 1/40 mailing labels per min
Second computer prints at the rate of 1/60 mailing labels per min
So:
1/40 + 1/60 = 1/x multiply each term by 120x
3x+2x=120
5x=120
x=24 min---------time required when both computers work together
CK
In 24 min first computer completes (1/40)*24=24/40 of the mailing labels
In 24 min second computer completes (1/60)*24=24/60 of the mailing labels
So 24/60 +24/40 should equal 1 (job that is)
24/60=48/120
24/40=72/120
48/120 + 72/120 = 120/120=1 job
Hope this helps--ptaylor